OpenStudy (lilah):

How would the expression x^3-3square root of 3 be rewritten using difference of cubes? A. (x+square root of 3)(x^2+square root of 3 +3) B. (x-square root of 3)(x^2+3x+3) C. (x-square root of 3)(x^2-3x-3) D. (x-square root of 3)(x^2+square root of 3x+3)

OpenStudy (lilah):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Is the original expression this? \[\Large x^3 - 3\sqrt{3}\] or is it this? \[\Large x^3 - \sqrt[3]{3}\]

OpenStudy (lilah):

the first one

jimthompson5910 (jim_thompson5910):

ok let me think

jimthompson5910 (jim_thompson5910):

notice how if \[\Large y = \sqrt{3}\] then cubing both sides gives \[\Large y^3 = (\sqrt{3})^3\] \[\Large y^3 = \sqrt{3}*\sqrt{3}*\sqrt{3}\] \[\Large y^3 = \sqrt{3*3*3}\] \[\Large y^3 = \sqrt{9*3}\] \[\Large y^3 = \sqrt{9}*\sqrt{3}\] \[\Large y^3 = 3\sqrt{3}\]

jimthompson5910 (jim_thompson5910):

So \[\Large x^3 - 3\sqrt{3}\] is in the form \[\Large x^3 - y^3\] where \[\Large y = \sqrt{3}\]

jimthompson5910 (jim_thompson5910):

From here, use this formula \[\Large x^3-y^3 = (x-y)(x^2+xy+y^2)\]

OpenStudy (lilah):

so i plug in 3 where the ys are?

jimthompson5910 (jim_thompson5910):

square root of 3

OpenStudy (lilah):

so i just put 1.7 where the ys are right?

jimthompson5910 (jim_thompson5910):

leave it in radical form

jimthompson5910 (jim_thompson5910):

your answer choices have sqrt(3) in them (not 1.7)

OpenStudy (lilah):

ohh right x) lol ok

OpenStudy (lilah):

thanks

jimthompson5910 (jim_thompson5910):

no problem